OpenStudy (joshoyen):

Use the properties of logarithms and the values below to find the logarithm indicated. Do not use a calculator to evaluate the logs. 1. \[\log_{8}6 \approx0.9\] \[\log_{8}5\approx0.8 \] \[\log_{8}9\approx1.1 \] Find \[\log_{8} \frac{ 1 }{ 36 }\]

OpenStudy (joshoyen):

Use the properties of logarithms and the values below to find the logarithm indicated. Do not use a calculator to evaluate the logs. 1. \[\log_{8}6 \approx0.9\] \[\log_{8}5\approx0.8 \] \[\log_{8}9\approx1.1 \] Find \[\log_{8} \frac{ 1 }{ 36 }\]

OpenStudy (campbell_st):

well take the question are rewrite it \[\log_{8} \frac{1}{36} = \log_{8}\frac{1}{6^2}\] the log laws for division says subtract the logs \[\log_{8}\frac{1}{6^2} = \log_{8}(1) - \log_{8}(6^2)\] you now need to use the power law for logs \[\log_{8}(1) - 2 \times \log_{8} (6)\] just evaluate it... and I hope it helps

OpenStudy (joshoyen):

Thanks!